The International Space Station, which has a mass of 4.96 x 10^5 kg, orbits 256 miles above the Earth's surface and completes one orbit every 88.0 minutes. What is the kinetic energy of the International Space Station in units of GJ (10^9 ​9 ​​ Joules)? (Note: don't forget to take into account the radius of the Earth!)
Answer #1

velocity of satellite

v = 2 pi r / t

v = 2 pi* ( 6.38* 10^6 + 4.12* 10^5) / (88* 60)

v = 8078.36 m/s

kE of energy

kE = 0.5* m v^2 = 0.5* 4.96 * 10^5* 8078.36^2

kE = 16184 GJ

==========

comment before rate in case any doubt, will reply for sure.. goodluck